# Set and functional prediction: randomness, exchangeability, and conformal

### v1

@@ 3.1 Kolmogorov’s step
3.2 Training-invariance step
@@
$\sigma$ ranging over the permutations of $\{1,\dots,n\}$. Such $G$ form a subclass of $\mathcal{E}^{\mathrm{X}}$.
@@
The interpretation of (11) is similar to that of (9).
@@
It suffices to check that the right-hand side of (11) is a test-conditional exchangeability e-variable. We have:
@@
3.3 Putting everything together
The following theorem combines Theorems 3 and 6 and establishes a connection between randomness and conformal e-predictors. Remember that the conformal e-predictor $\mathcal{E}^{\mathrm{tX}}$ derived from a randomness e-predictor $E$ is obtained by combining the operators (6) and (7), i.e., as
$E^{\mathrm{tX}}(z_{1},\dots,z_{n+1}):=(n+1)\frac{\sum_{\sigma}E(z_{\sigma(1)},\dots,z_{\sigma(n)},z_{n+1})}{\sum_{\pi}E(z_{\pi(1)},\dots,z_{\pi(n+1)})},$ (12)
@@ 3.3 Putting everything together
Theorem 7 is the main result of this paper for e-predictors. Its main weakness is the presence of the term $\mathrm{e}^{-1/2}$, but it might be inevitable.
@@
Applying the Cauchy–Schwarz inequality, we have, for some $G_{1},G_{2},G\in\mathcal{E}^{\mathrm{R}}$,
$\mathrm{e}^{-1/2}\int\sqrt{\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},y)}}\;B(\mathrm{d}y\mid z_{n+1})\\ =\mathrm{e}^{-1/2}\int\sqrt{\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}}\sqrt{\frac{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},y)}}\;B(\mathrm{d}y\mid z_{n+1})\\ \leq\sqrt{\mathrm{e}^{-1}\int\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}\;B(\mathrm{d}y\mid z_{n+1})}\\ \quad{}\times\sqrt{\int\frac{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},y)}B(\mathrm{d}y\mid z_{n+1})}\\ =\sqrt{G_{1}(z_{1},\dots,z_{n+1})G_{2}(z_{1},\dots,z_{n+1})}\leq G(z_{1},\dots,z_{n+1})$
(the existence of $G_{1}$ and $G_{2}$ follows from Theorems 3 and 6, respectively). ∎
It is known that, for any $\delta\in(0,1)$, the function $p\mapsto\delta p^{\delta-1}$ transforms p-values to e-values and that the function $e\mapsto e^{-1}$ transforms e-values to p-values (see, e.g., [11, Propositions 2.1 and 2,2]). This allows us to adapt Theorem 8 to p-predictors.
Theorem 8.
@@
$G(z_{1},\dots,z_{n},z_{n+1}):=(\delta/\mathrm{e})^{1/2}\\ \times\int\sqrt{P(z_{1},\dots,z_{n},x_{n+1},y)^{\delta-1}P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)}\>B(\mathrm{d}y\mid z_{n+1})$ (14)
@@
The interpretation of (14) is that $P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)$ is typically small (perhaps not to the same degree) when $P(z_{1},\dots,z_{n},x_{n+1},y)$ is small; i.e., we do not lose much in efficiency when converting randomness p-predictors to conformal p-predictors. To see this, fix small $\epsilon_{1},\epsilon_{2}\in(0,1)$. Then we will have $G(z_{1},\dots,z_{n},z_{n+1})<1/\epsilon_{1}$ for the true data sequence $z_{1},\dots,z_{n},z_{n+1}$ unless a rare event (of probability at most $\epsilon_{1}$) happens. For the vast majority of the potential labels $y\in\mathbf{Y}$ we will have
$(\delta/\mathrm{e})^{1/2}\sqrt{P(z_{1},\dots,z_{n},x_{n+1},y)^{\delta-1}P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)}<\frac{1}{\epsilon_{1}\epsilon_{2}},$ (15)
$P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)<\frac{\mathrm{e}P(z_{1},\dots,z_{n},x_{n+1},y)^{1-\delta}}{\delta\epsilon_{1}^{2}\epsilon_{2}^{2}},$
Proof of Theorem 8.
Fix $\delta\in(0,1)$ and $P\in\mathcal{P}^{\mathrm{R}}$. Set $E:=\delta P^{\delta-1}$ and $P^{\prime}:=1/E^{\mathrm{tX}}$, so that $E\in\mathcal{E}^{\mathrm{R}}$ and $P^{\prime}\in\mathcal{P}^{\mathrm{tX}}$. According to (13),
$G(z_{1},\dots,z_{n},z_{n+1}):=\mathrm{e}^{-1/2}\int\sqrt{\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},y)}}\;B(\mathrm{d}y\mid z_{n+1})\\ =\mathrm{e}^{-1/2}\int\sqrt{\delta P(z_{1},\dots,z_{n},x_{n+1},y)^{\delta-1}P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)}B(\mathrm{d}y\mid z_{n+1})$
is a randomness e-variable. ∎
4 Applications to classification
In this and next sections we will discuss some interesting examples of Markov kernels $B$ as used in Theorems 7 and 8. In this section we discuss the case of classification, $\left|\mathbf{Y}\right|<\infty$, which was also discussed earlier in [8].
Let us start from binary classification, $\mathbf{Y}:=\{0,1\}$. In this case, the most natural choice of $B$ is $B(\{y\}\mid(x,y)):=0$, so that the Markov kernel sends every example $(x,y)$ to the other label $1-y$. We can rewrite (13) as
$G(z_{1},\dots,z_{n},z_{n+1}):=\mathrm{e}^{-1/2}\sqrt{\frac{E(z_{1},\dots,z_{n},x_{n+1},1-y_{n+1})}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},1-y_{n+1})}},$ (16)
which does not involve any averaging. We can interpret (16) as the conformal e-predictor $E^{\mathrm{tX}}$ being almost as efficient as the original randomness e-predictor $E$, where efficiency is measured by the degree to which we reject the false label $1-y_{n+1}$. For example, for a small positive constant $\epsilon$, $G\geq 1/\epsilon$ with probability at most $\epsilon$, and so
$E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},1-y_{n+1})>\mathrm{e}^{-1}\epsilon^{2}E(z_{1},\dots,z_{n},x_{n+1},1-y_{n+1})$
with probability at least $1-\epsilon$.
In the case of reduction of a randomness p-predictor, we rewrite (14) as
$G(z_{1},\dots,z_{n},z_{n+1})=(\delta/\mathrm{e})^{1/2}\\ {}\times\sqrt{P(z_{1},\dots,z_{n},x_{n+1},1-y_{n+1})^{\delta-1}P^{\prime}(z_{1},\dots,z_{n},x_{n+1},1-y_{n+1})}.$
Therefore,
$P^{\prime}(z_{1},\dots,z_{n},x_{n+1},1-y_{n+1})<\mathrm{e}\delta^{-1}\epsilon^{-2}P(z_{1},\dots,z_{n},x_{n+1},1-y_{n+1})^{1-\delta}$
with probability at least $1-\epsilon$. The interpretation is similar to that of the e-version.
In the rest of this section, let us only discuss reduction of randomness e-predictors to conformal e-predictors. Reduction of randomness p-predictors to conformal p-predictors is completely analogous; it just uses (14) instead of (13).
$G(z_{1},\dots,z_{n},z_{n+1}):=\frac{\mathrm{e}^{-1/2}}{\left|\mathbf{Y}\right|-1}\sum_{y\in\mathbf{Y}\setminus\{y_{n+1}\}}\sqrt{\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},y)}}.$ (17)
The interpretation of (17) is that the conformal e-predictor $E^{\mathrm{tX}}$ is almost as efficient as the original randomness e-predictor $E$ on average; as before, efficiency is measured by the degree to which we reject the false labels $y\neq y_{n+1}$. Of course, we can avoid “on average” by making (17) cruder and replacing it by the existence of $G\in\mathcal{E}^{\mathrm{R}}$ satisfying
$\forall(z_{1},\dots,z_{n})\in\mathbf{Z}^{n}\;\forall x_{n+1}\in\mathbf{X}\;\forall y\in\mathbf{Y}\setminus\{y_{n+1}\}:\\ G(z_{1},\dots,z_{n},z_{n+1})\geq\frac{\mathrm{e}^{-1/2}}{\left|\mathbf{Y}\right|-1}\sqrt{\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},y)}},$
where $z_{n+1}:=(x_{n+1},y_{n+1})$. For a small positive constant $\epsilon$, we can then claim that, with probability at least $1-\epsilon$,

### v2

@@ A.2 Proof of Theorem 6
A.3 Proof of Corollary 7
Fix $\delta\in(0,1)$ and $P\in\mathcal{P}^{\mathrm{R}}$. Set $E:=\delta P^{\delta-1}$ and $P^{\prime}:=1/E^{\mathrm{X}}$, so that $E\in\mathcal{E}^{\mathrm{R}}$ and $P^{\prime}\in\mathcal{P}^{\mathrm{X}}$. According to (7),
$G(z_{1},\dots,z_{n},z_{n+1})$ $:=\mathrm{e}^{-1}\int\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}\;B(\mathrm{d}y\mid z_{n+1})$
$=\mathrm{e}^{-1}\int\frac{\delta P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)}{P^{1-\delta}(z_{1},\dots,z_{n},x_{n+1},y)}\;B(\mathrm{d}y\mid z_{n+1})$
is a randomness e-variable.
A.4 Proof of Theorem 8
The examples $z_{1},\dots,z_{n+1}$ whose existence is asserted in the statement of the theorem and which will be constructed in this proof will all share the same fixed object $x_{0}\in\mathbf{X}$, and we will sometimes omit “$x_{0}$” in our notation. Suppose, without loss of generality, that the label set $\mathbf{Y}$ is the disjoint union of $\{0,1\}$ and the set $\{-k,\dots,k\}$ for some positive integer $k$; to distinguish between the 0s and the 1s in these two disjoint sets, let us write $0^{\prime}$ and $1^{\prime}$ for the elements of the first set, $\{0,1\}=\{0^{\prime},1^{\prime}\}$. (The primes are ignored, of course, when $0^{\prime}$ and $1^{\prime}$ are used as inputs to arithmetic operations, as in (25) below. If $\left|\mathbf{Y}\right|$ is an even number, we can leave one of its elements unused.) To avoid trivial complications, let $n$ be an even number. This proof will assume $1\ll k\ll\sqrt{n}$; the formal meaning of this assumption will be summarized at the end of the proof.
Define $E\in\mathcal{E}^{\mathrm{tR}}$ (which ignores the objects) as follows:
  • on the sequences in $\mathbf{Z}^{n+1}$ of the form $(x_{1},y_{1},\dots,x_{n},y_{n},x_{n+1},y)$, where $y_{1},\dots,y_{n}\in\{0^{\prime},1^{\prime}\}$, $y\in\{-k,\dots,k\}$, and
    $\sum_{i=1}^{n}y_{i}-n/2=y,$ (25)
    $E$ takes value $a\mathrm{e}\sqrt{\pi/2}n^{3/2}$, where $a<1$ is a positive constant (it will be taken close to 1 later in the proof);
  • $E$ takes value 0 on all other sequences in $\mathbf{Z}^{n+1}$.
In (25), $y$ is determined by $y_{1}+\dots+y_{n}$ and vice versa. This agrees with the idea of the proof of Theorem 8 given after its statement: since $y_{1},\dots,y_{n}$ are binary, $y_{1}+\dots+y_{n}$ carries the same information as $\lbag y_{1},\dots,y_{n}\rbag$.
Let us check that $E$ is indeed a train-invariant randomness e-variable. Let the underlying probability space be $\mathbf{Z}^{n+1}$ equipped with a probability measure $R=Q^{n+1}$, so that individual examples are generated independently from $Q$. We will use the notation $Z_{i}$, $i=1,\dots,n+1$, for $z_{i}$ considered as a random example (formally, $Z_{i}$ is the random element defined by $Z_{i}(z_{1},\dots,z_{n+1}):=z_{i}$) and the notation $Y_{i}$, $i=1,\dots,n+1$, for the label of $Z_{i}$. The maximum probability of the event $E(Z_{1},\dots,Z_{n+1})>0$ is attained for $Q$ giving maximum probabilities to the following two events:
  1. 1.
    The random labels $Y_{1},\dots,Y_{n}$ take values in $\{0^{\prime},1^{\prime}\}$, and the remaining random label $Y_{n+1}$ takes value in $\{-k,\dots,k\}$.
  2. 2.
    Conditionally on the first event, we have (25), where $y$ is the value taken by $Y_{n+1}$ and $y_{1},\dots,y_{n}$ are the values taken by $Y_{1},\dots,Y_{n}$.
The maximum probability of the first event (in item 1) is
$\left(1-\frac{1}{n+1}\right)^{n}\frac{1}{n+1}\sim\frac{1}{\mathrm{e}n},$
which is obtained by maximizing $(1-\theta)^{n}\theta$ over $\theta$ and where the asymptotic equivalence is as $n\to\infty$. The maximum probability of the second event (in item 2) conditional on both the first event and $Y_{n+1}=y\in\{-k,\dots,k\}$ is asymptotically equivalent, by the local limit theorem [12, Sect. 1.6], to
$\frac{1}{\sqrt{2\pi n\left(\frac{1}{2}+\frac{y}{n}\right)\left(\frac{1}{2}-\frac{y}{n}\right)}}=\frac{1}{\sqrt{2\pi n\left(\frac{1}{4}-\frac{y^{2}}{n^{2}}\right)}}\sim\sqrt{\frac{2}{\pi n}};$
this follows from the random variables $Y_{1},\dots,Y_{n}$ being distributed as $B_{\theta}^{n}$, and the maximum over $\theta$ being attained at $\theta=1/2+y/n$. Since $a<1$, $E\in\mathcal{E}^{\mathrm{tR}}$ for a sufficiently large $n$. Notice that our argument in this paragraph only uses $k\ll n$, since in our application of the local limit theorem the exponential term $\exp(\dots)$ was 1.
Let us now generate $Y_{1},\dots,Y_{n+1}$ from the Bernoulli model and find the maximum probability that
$\sum_{i=1}^{n}Y_{i}-n/2\in\{-k,\dots,k\}$ (26)
(so that this condition does not involve $Z_{n+1}$). Again by the local limit theorem, the maximum probability is asymptotically equivalent to
$\frac{2k+1}{\sqrt{2\pi n/4}}\sim\frac{2\sqrt{2}k}{\sqrt{\pi n}};$
it is attained at $\theta:=1/2$. Now the assumption $1\ll k\ll\sqrt{n}$ is essential in order for the exponential term in the local limit theorem to go away. Therefore, for any $G\in\mathcal{E}^{\mathrm{R}}$, there are $y_{1},\dots,y_{n+1}\in\{0^{\prime},1^{\prime}\}$ such that
$G(y_{1},\dots,y_{n+1})\leq\frac{b\sqrt{\pi n}}{2\sqrt{2}k}$
(remember that we are omitting $x_{0}$), where $b>1$ is a constant (to be chosen close to $1$ later), and
$\sum_{i=1}^{n}y_{i}-n/2\in\{-k,\dots,k\}$
(cf. (26)). Fix such $y_{1},\dots,y_{n+1}$ and set
$y:=\sum_{i=1}^{n}y_{i}-n/2$
(cf. (25)). Taking $a$ and $b$ sufficiently close to 1, we obtain
$\frac{G(y_{1},\dots,y_{n+1})}{E^{\mathrm{i}}(y_{1},\dots,y_{n},y)}=(n+1)\frac{G(y_{1},\dots,y_{n+1})}{E(y_{1},\dots,y_{n},y)}\leq\frac{b\sqrt{\pi n}(n+1)}{2\sqrt{2}ka\mathrm{e}\sqrt{\pi/2}n^{3/2}}<\frac{c}{\mathrm{e}\left|\mathbf{Y}\right|}$ (27)
for a sufficiently large $k$. The equality “$=$” in (27) follows from $\{0^{\prime},1^{\prime}\}$ and $\{-k,\dots,k\}$ being disjoint sets, the inequality $\leq$” follows from the definitions of $E$ and $G$, and the inequality “$<$” follows from $\left|\mathbf{Y}\right|=2k+3$. This proves Theorem 8 since $E^{\mathrm{i}}=E/E^{\mathrm{X}}$.
Finally, the formal meaning of the condition $1\ll k\ll\sqrt{n}$ is that the first sentence in the statement of Theorem 8 can be replaced by “For each constant $c>1$ there is $C>0$ such that the following statement holds true assuming $\left|\mathbf{Y}\right|\geq C$ and $\sqrt{n}\geq C\left|\mathbf{Y}\right|$.”
A.5 Proof of Theorem 9
This section spells out details omitted in the proof sketch given in Sect. 4. As mentioned there, the probability of (20) being at least $0.5$ and the probability of (21) being at least $0.502$ follow immediately from Markov’s inequality. The notation $Y_{1},\dots,Y_{n+1}$ was introduced formally in Sect. A.4 (and used informally already in the proof sketch of Theorem 9).
Now let us define $E$: {wideformula} E(y_1,…,y_n+1) :=
{c—Yif $S\equiv 0\ (\mathrm{mod}\,\left|\mathbf{Y}\right|)$ and $\forall y\in\mathbf{Y}:\bigl{|}k_{y}-n/\left|\mathbf{Y}\right|\bigr{|}\leq 0.1n/\left|\mathbf{Y}\right|$0 otherwise, where $S$ is the sum of $y_{1},\dots,y_{n+1}$ and $k_{y}:=\left|\left\{i\in\{1,\dots,n+1\}\mid y_{i}=y\right\}\right|$ is the number of times $y\in\mathbf{Y}$ occurs in the sequence $y_{1},\dots,y_{n+1}$. We are required to prove two statements for a large enough $n$: first, that $E\in\mathcal{E}^{\mathrm{R}}$, and second, that the probability of $E(Y_{1},\dots,Y_{n},Y)>0$ is at least 0.999 (see (22)).
The second statement is easier. Let $y_{1},\dots,y_{n+1}$ be the values taken by $Y_{1},\dots,Y_{n},Y$, respectively. Then the first condition
$S\equiv 0\ (\mathrm{mod}\,\left|\mathbf{Y}\right|)$ (28)
in (A.5) holds automatically, and the second condition
$\forall y\in\mathbf{Y}:\bigl{|}k_{y}-n/\left|\mathbf{Y}\right|\bigr{|}\leq 0.1n/\left|\mathbf{Y}\right|$ (29)
holds with probability that tends to 1 as $n\to\infty$ for each $y\in\mathbf{Y}$, by the central limit theorem. Since $\mathbf{Y}$ is a finite set, the second statement has been proved.
Now let us prove the first statement. Let $Q$ be a probability measure on $\mathbf{Y}$ and set $R:=Q^{n+1}$. We will denote by $A$ the event given after “if” in (A.5), namely the conjunction of (28) and (29). It suffices to prove that asymptotically as $n\to\infty$ the $R$-probability of $A$ does not exceed $c^{-1/2}/\left|\mathbf{Y}\right|$, uniformly in $Q$. Consider two cases:
  • If $Q(\{y\})\geq 0.1/\left|\mathbf{Y}\right|$ for all $y\in\mathbf{Y}$, the convergence $R(A_{1})\to 1/\left|\mathbf{Y}\right|$ as $n\to\infty$ for the superset $A_{1}$ of $A$ given by (28) follows from the asymptotic uniformity result of [7, Theorem (5.01)] (proved independently in [3, Theorem 2]). The convergence is uniform on this compact set of $Q$.
  • If $Q(\{y\})\leq 0.1/\left|\mathbf{Y}\right|$ for some $y\in\mathbf{Y}$, the convergence $R(A_{2})\to 0$ as $n\to\infty$ for the superset $A_{2}$ of $A$ given by (29) follows from, e.g., the central limit theorem combined with the Bonferroni correction for the multiplicity of $y$. The convergence is again uniform on this compact set of $Q$.
Combining the two cases, we get $E\in\mathcal{E}^{\mathrm{R}}$.
Appendix B Reducing randomness predictors to conformal
In the main paper we assumed the train-invariance of randomness predictors. In this appendix we drop this assumption, since statistical principles may go astray (see, e.g., the criticism [14] of the conditionality principle from the vantage point of conformal prediction). Besides, in some cases we may have convincing reasons to violate train-invariance, such as using inductive conformal predictors [17, Sect. 4.2] for the purpose of computational efficiency.
[…]
Figure 2: A cube representing the main classes of predictors considered in Appendix B. The right panel shows embeddings as arrows, as in Figure 1.
The connections between various classes of predictors (introduced in the main paper) that we use or explore in this appendix are shown in red in the left-hand panel of Figure 2. Namely, we are interested in connections between:
  1. (a)
    randomness e-predictors and exchangeability e-predictors;
  2. (b)
    exchangeability e-predictors and conformal e-predictors;
  3. (c)
    randomness p-predictors and randomness e-predictors;
  4. (d)
    conformal e-predictors and conformal p-predictors.
Theorem 5 stated in Subsect. 3.1 is general enough to cover connection (a). Connection (b) will be discussed in Sect. B.2. Connections (c) and (d) result from the possibility of converting p-values to e-values and back, as discussed in Subsect. 3.2.
In Sect. B.3 we will see how the connections shown in red in Figure 2 can be combined to demonstrate the universality of conformal prediction without accepting the train-invariance principle. Section B.4 applies this to classification. But we start in Sect. B.1 by introducing an operator making randomness predictors (in particular, exchangeability predictors) train-invariant.
B.1 Another operator
In addition to the operators (5) and (6), we will need another operator,
$\sigma$ ranging over the permutations of $\{1,\dots,n\}$. Now we have $E^{\mathrm{t}}\in\mathcal{E}^{\mathrm{tX}}$ (resp. $E^{\mathrm{t}}\in\mathcal{E}^{\mathrm{tR}}$) whenever $E\in\mathcal{E}^{\mathrm{X}}$ (resp. $E^{\mathrm{t}}\in\mathcal{E}^{\mathrm{R}}$). Both operators (5) and (30) are kinds of averaging: while $E\mapsto E^{\mathrm{i}}$ averages over all permutations of an input data sequence (including both training and test examples), $E\mapsto E^{\mathrm{t}}$ averages over the permutations of the training sequence only. And while $E\mapsto E^{\mathrm{i}}$ is polynomially computable for train-invariant $E$, $E\mapsto E^{\mathrm{t}}$ requires exponential computation time in general.
Using two of the three operators (5), (6), and (30), we can turn any randomness e-variable $E$ to an exchangeability e-variable $E^{\mathrm{X}}$ to a conformal e-variable $(E^{\mathrm{X}})^{\mathrm{t}}$. The following lemma shows that the order in which the last two operators are applied does not matter.
Lemma 11.
Proof.
Let us fix a data sequence $z_{1},\dots,z_{n+1}$ and check
$(E^{\mathrm{t}})^{\mathrm{X}}(z_{1},\dots,z_{n+1})=(E^{\mathrm{X}})^{\mathrm{t}}(z_{1},\dots,z_{n+1}).$
As functions of a permutation of $z_{1},\dots,z_{n+1}$, $E$ and $E^{\mathrm{X}}$ are proportional to each other, and therefore, $E^{\mathrm{t}}$ and $(E^{\mathrm{X}})^{\mathrm{t}}$ are also proportional to each other. This implies $(E^{\mathrm{t}})^{\mathrm{X}}=(E^{\mathrm{X}})^{\mathrm{t}}$ on the permutations of $z_{1},\dots,z_{n+1}$. And this is true for each $(z_{1},\dots,z_{n+1})$. ∎
We will let tX stand for the composition of the two operators:
$E^{\mathrm{tX}}:=(E^{\mathrm{t}})^{\mathrm{X}}=(E^{\mathrm{X}})^{\mathrm{t}}.$
Let us say that $E\in\mathcal{E}^{\mathrm{X}}$ is admissible if (4) always holds with “$=$” in place of $\leq$”. (This agrees with the standard notion of admissibility in statistical decision theory.) The intuition (which can be formalized easily) behind the four operators that we have introduced in this paper is that:
  • t projects $\mathcal{E}^{\mathrm{X}}$ onto $\mathcal{E}^{\mathrm{tX}}$ and $\mathcal{E}^{\mathrm{R}}$ onto $\mathcal{E}^{\mathrm{tR}}$;
In particular, these operators are idempotent:
$(E^{\mathrm{i}})^{\mathrm{i}}=E^{\mathrm{i}},\quad(E^{\mathrm{X}})^{\mathrm{X}}=E^{\mathrm{X}},\quad(E^{\mathrm{t}})^{\mathrm{t}}=E^{\mathrm{t}},\quad(E^{\mathrm{tX}})^{\mathrm{tX}}=E^{\mathrm{tX}}$
(and we can allow any $E\in\mathcal{E}$ here). Despite these operators being projections, we cannot claim that these ways of moving between different classes of predictors are always optimal.
B.2 Efficiency of making predictors train-invariant
@@
$\sigma$ ranging over the permutations of $\{1,\dots,n\}$. Such $G$ form a subclass of $\mathcal{E}^{\mathrm{X}}$ (and therefore, of $\mathcal{E}^{\mathrm{R}}$).
@@
The interpretation of (31) is similar to that of (7). It will be clear from the proof that we can allow $E$ to be any randomness e-predictor, or even any element of $\mathcal{E}$.
@@
Let us check that the right-hand side of (31) is a test-conditional exchangeability e-variable:
@@
B.3 Putting everything together
The following theorem combines Theorems 5 and 12 and establishes a connection between randomness and conformal e-predictors, without accepting the train-invariance principle. Remember that the conformal e-predictor $\mathcal{E}^{\mathrm{tX}}$ derived from a randomness e-predictor $E$ is obtained by combining the operators (6) and (30), i.e., as
$E^{\mathrm{tX}}(z_{1},\dots,z_{n+1}):=(n+1)\frac{\sum_{\sigma}E(z_{\sigma(1)},\dots,z_{\sigma(n)},z_{n+1})}{\sum_{\pi}E(z_{\pi(1)},\dots,z_{\pi(n+1)})},$
@@ B.3 Putting everything together
The main weakness of Corollary 13 is the presence of the term $\mathrm{e}^{-1/2}$, but it might be inevitable.
@@
Applying the Cauchy–Schwarz inequality, we have, for some $G_{1},G_{2},G_{3}\in\mathcal{E}^{\mathrm{R}}$,
$\mathrm{e}^{-1/2}\int\sqrt{\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},y)}}\;B(\mathrm{d}y\mid z_{n+1})\\ =\mathrm{e}^{-1/2}\int\sqrt{\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}}\enspace\sqrt{\frac{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},y)}}\;B(\mathrm{d}y\mid z_{n+1})\\ \leq\sqrt{\mathrm{e}^{-1}\int\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}\;B(\mathrm{d}y\mid z_{n+1})}\\ \quad{}\times\sqrt{\int\frac{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},y)}B(\mathrm{d}y\mid z_{n+1})}\\ =\sqrt{G_{1}(z_{1},\dots,z_{n+1})G_{2}(z_{1},\dots,z_{n+1})}\leq G_{3}(z_{1},\dots,z_{n+1})$
(the existence of $G_{1}$ and $G_{2}$ follows from Theorems 5 and 12, respectively, and $G_{3}$ can be set, e.g., to the arithmetic average of $G_{1}$ and $G_{2}$). ∎
Similarly to Corollary 7, we can adapt Corollary 13 to p-predictors.
Corollary 14.
@@
$G(z_{1},\dots,z_{n},z_{n+1}):=\sqrt{\frac{\delta}{\mathrm{e}}}\int\sqrt{\frac{P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)}{P^{1-\delta}(z_{1},\dots,z_{n},x_{n+1},y)}}\>B(\mathrm{d}y\mid z_{n+1})$ (33)
@@
The interpretation of (33) is similar to that of (10): $P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)$ is typically small when $P(z_{1},\dots,z_{n},x_{n+1},y)$ is small. Instead of (11) we will have
$\sqrt{\frac{\delta}{\mathrm{e}}}\sqrt{\frac{P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)}{P^{1-\delta}(z_{1},\dots,z_{n},x_{n+1},y)}}<\frac{1}{\epsilon_{1}\epsilon_{2}},$
which can be rewritten as
$P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)<\frac{\mathrm{e}}{\delta\epsilon_{1}^{2}\epsilon_{2}^{2}}P^{1-\delta}(z_{1},\dots,z_{n},x_{n+1},y)$
in place of (12).
Proof of Corollary 14.
We proceed as in the proof of Corollary 7 except for replacing (7) by (32). ∎
B.4 Applications to classification
As compared with Sect. 4, we get similar but weaker performance guarantees for the derived conformal predictors without accepting the train-invariance principle. In the binary case $\mathbf{Y}=\{-1,1\}$, we apply (32) to obtain
$G(z_{1},\dots,z_{n},z_{n+1}):=\mathrm{e}^{-1/2}\sqrt{\frac{E(z_{1},\dots,z_{n},x_{n+1},-y_{n+1})}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},-y_{n+1})}}$
in place of (8). This implies (9) with $\mathcal{E}^{\mathrm{tX}}$ in place of $\mathcal{E}^{\mathrm{X}}$ and $\epsilon^{2}$ in place of $\epsilon$.
Instead of (13) now we have
$G(z_{1},\dots,z_{n},z_{n+1}):=\frac{\mathrm{e}^{-1/2}}{\left|\mathbf{Y}\right|-1}\sum_{y\in\mathbf{Y}\setminus\{y_{n+1}\}}\sqrt{\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{tX}}(z_{1},\dots,z_{n},x_{n+1},y)}},$
and instead of (15) we have